The graph of the function is shown. Find (or estimate) each of the items below. If the symbol oo oro applies, use it. If a li A - H
Answer #1

b) solution Al to graph. Ô f(1) = ? solution :- at nil, then value of y = -3( In graph is denoted by dot) . Therefore value T(C) lim fu fon)] = ? (2) page. ->29 2 Now f(n) from graph. ? when a 1 +4E ........ then slople of line (m) f(a) f il 4 -0 -+@ $112) = ? 3) page The geometrical meaning of f(n) mean slope of line of y = f(u) at u=2. In previous question C? we had fiу. (Чј рок - () [460-1 )du = ? Azou a y equation of line 1 from Jan Area = 8 fuldre.o how slope of above line from point (0,1may be helpful for you...if you have any quires please discuss with me ,, I will help you...!!